OpenStudy (anonymous):

Solve by factoring: x^2+3x=18

OpenStudy (anonymous):

x^2+3x-18=0 (x+9)(x-6)=0 x=-9,6

OpenStudy (anonymous):

that wrong the factors of 18=1,2,3,6,9,19 correct answer (x-3)(x+6)=0 x=3 , x=-6

OpenStudy (anonymous):

the factor pairs of 18 are 1,18 2,9 3,6